OpenStudy (anonymous):

Producing a musical costs $88,00 plus $5600 per performance. One sold-out performance earns $7500 in revenue. If every performance sells out, how many performances are needed to break even?

OpenStudy (anonymous):

If we take this work problem we can easily make an equation out of it So Let us say Y is how much money we make Y=-8800+5600x+7500z x is the number of performances z is the number of sold out performances I think you may have missed a 0 in 88000? if so you can just add the extra 0. From there, you will want to set Y=0, which means they did not lose any money or gain any money, breaking even. and with only sell out performances you can set x=0, so our new equation is Y=-8800+7500z 0=-8800+7500z solve for z :)

OpenStudy (anonymous):

So sorry, it should be: Y=-8800-5600x+7500x x is your number of performances :)

OpenStudy (anonymous):

So just to clear it up one step more You can combine the two x terms so... Y=-8800+1900x Plug 0 in for Y and solve for x :)

OpenStudy (anonymous):

Thanks so much :)